Find a recurrence relation, with initial condition, that uniquely determines each of the following geometric progressions.

a)        2, 10, 50, 250, . . .


                

b)        7, 14/5, 28/25, 56/125, . . .

 


1
Expert's answer
2021-11-18T07:00:37-0500

 Given Geometric progression;

a) 2,10,50,250,

Rewrite as,

"\begin{aligned}

&a_{0}=2 \\

&a_{1}=10 \\

&a_{2}=50 \\

&a_{3}=250

\end{aligned}"

 Now change each term of sequence in terms of a_{0};

 "\begin{aligned}

\text { since } a_{1} &=10=5 \times 2 \\

& \Rightarrow a_{1}=5 a_{0} \\

a_{2}=50 &=25 \times 2=25 a_{0} \\

& \Rightarrow a_{2}=(5)^{2} a_{0} \\

 a_{3}=250 &=125 \times 2 \\

&=(5)^{3} a_{0}

\end{aligned}"

Therefore, the recurrence relation is given by

"a_{n}=(5)^{n} a_{0} \quad \forall n \geqslant 1"

with Initial Condition "a_{0}=2"


b)

"\begin{array}{ll}7, 14 / 5,28 / 25,56 / 125, \ldots \\ a_{0}=7, a_{1}=14/5, a_{2}=28/25, a_{3}=56/125\end{array}"

Now,

"\begin{aligned}

a_{1}=\frac{14}{5} &=\frac{2 \times 7}{5} \\

\Rightarrow a_{1} &=\frac{2}{5} a_{0}

\end{aligned}"

 

 "\begin{aligned}

a_{2}=\frac{28}{25} &=\frac{4}{25} \times 7=\left(\frac{2}{5}\right)^{2} \times 7 \\

& \Rightarrow a_{2}=\left(\frac{2}{5}\right)^{2} a_{0} \mid

\end{aligned}"

 

"\begin{aligned}

a_{3}=\frac{56}{125} &=\frac{8 \times 7}{125}=\left(\frac{2}{5}\right)^{3} \times 7 \\

& \Rightarrow a_{3}=\left(\frac{2}{5}\right)^{3} a_{0}

\end{aligned}"

 Therefore Recurrence relation is given by;

"\ a_{n}=\underbrace{\left(\frac{2}{5}\right)^{n} a_{0} \quad \forall n \geqslant 1}_{\text {with } a_{0}=7}"

OR

"a_{n}=\frac{2}{5} a_{n-1}"